# StreamDiffusionV2: A Streaming System for Dynamic and Interactive Video Generation

### v1

@@ 3 Motivation & Bottleneck Analysis
Figure 4: Roofline analysis of sequence parallelism and our pipeline orchestration. We compare the Sequence Parallelism and Pipeline Parallelism under varying batch sizes in the causal DiT, compared with the bidirectional DiT. The results demonstrate that our approach operates near the knee point of the roofline, effectively avoiding compute under-utilization as seen in the bidirectional DiT and memory bandwidth limitations in Sequence Parallelism. The model is profiled on an NVIDIA H100 SXM GPU with a peak performance of 1,979 TFLOP/s and a knee point at an arithmetic intensity (AI) of 590.75 FLOP/Byte. The token length is 1,536 (a 4-frame chunk at 480P resolution.)
@@
In live stream scenarios, strict per-frame deadlines hinder the scalability of conventional parallelization strategies for two key reasons: (i) communication latency in sequence parallelism significantly reduces potential speedup, and (ii) short-frame chunks drive the workload into a memory-bound regime, as shown in Fig. 4. These effects are further amplified in real-time streaming, where efficient causal DiTs operate on short sequences (e.g., 4 frames per step), reducing per-frame computation and making communication overhead proportionally heavier (see Fig. 5).
@@ 5.3 Generation Quality Evaluation
CLIP Score $\uparrow$ 95.24 96.58 98.48 98.51
@@ 5.4.2 Effectiveness of the Dynamic DiT-Block Scheduler
To isolate algorithmic scaling, we evaluate the theoretical latency of sequence parallelism and pipeline parallelism with communication removed (Fig. 14). Our method, built on pipeline parallelism, attains a near-ideal acceleration by partitioning DiT blocks. In contrast, sequence parallelism shows clear gains only at high resolutions; at moderate and low resolutions, it shifts the workload into a memory-bound regime, yielding little to no latency reduction and underutilizing compute.
@@ 6 Conclusion
We propose StreamDiffusionV2, which closes the gap between offline video diffusion and live streaming constrained in real-time with SLO constraints. Our training-free system couples an SLO-aware batching/block scheduler with a sink-token–guided rolling KV cache, a motion-aware noise controller, and a pipeline orchestration that parallelizes across denoising steps and model layers—delivering near-linear FPS scaling without violating latency. It runs on heterogeneous GPUs and flexible step counts, achieving 0.5 s TTFF and up to 58.28 FPS (14B) / 64.52 FPS (1.3B) on 4×H100, and maintaining high FPS even as steps increase. These results make state-of-the-art generative live streaming practical for both individual creators and enterprise platforms.
[…]
7 NVIDIA H100 SXM Roofline Parameters
We derive the roofline parameters for the NVIDIA H100 SXM GPU. The dense FP16 throughput is $1,979\penalty 10000\ \mathrm{TFLOP/s}$ and GPU memory bandwidth is $3.35\penalty 10000\ \mathrm{TB/s}$, which are taken from the H100 datasheet  [^4] https://resources.nvidia.com/en-us-gpu-resources/h100-datasheet-24306. The ridge arithmetic intensity is
$\mathrm{AI}_{\mathrm{ridge}}=\frac{1979\penalty 10000\ \mathrm{TFLOP/s}}{3.35\penalty 10000\ \mathrm{TB/s}}=590.75\penalty 10000\ \mathrm{FLOP/Byte}$ (3)
8 REPA finetuning for quality enhancement
StreamDiffusionV2 is a training-free system solution to transform efficient video models for live streaming applications. Orthogonal to the training-free solution, we employ the REPA (yu2024repa) training strategy during Causal-DiT distillation (yin2025slow) to enhance the video quality. The alignment loss is denoted by
$\mathcal{L}_{\mathrm{REPA}}(\theta,\phi):=-\mathbb{E}_{\mathbf{x}_{*},\epsilon,t}\left[\frac{1}{N}\sum_{n=1}^{N}\mathrm{sim}\!\left(\mathbf{y}_{*}^{[n]},\,h_{\phi}\!\left(\mathbf{h}_{t}^{[n]}\right)\right)\right],$
where $\mathbf{y}_{*}^{[n]}$ indicates the output of DINOv2 (oquab2024dinov2), $\mathbf{h}_{t}^{[n]}$ represents the hidden state of DiT in timestep $t$, $h_{\phi}$ is the projection network parameterized by $\phi$, and $\mathrm{sim}(\cdot,\cdot)$ denotes cosine similarity. When combined with the original DMD (yin2024one) distillation objective $\mathcal{L}_{\mathrm{DMD}}$, the overall training objective is defined as
$\mathcal{L}=\mathcal{L}_{\mathrm{DMD}}+\lambda\mathcal{L}_{\mathrm{REPA}},$
where $\lambda$ is a scaling factor.
9 Figure illustration
Figure overview. Fig. 17 shows how we minimize pipeline bubbles by separating compute and communication into two concurrent streams, overlapping kernels with P2P transfers.
Fig. 17 illustrates the rolling KV cache with sink tokens for consistent, frame-by-frame updates.
Fig. 18 reports a comparison of parallelization strategies (e.g., sequence parallelism vs. our pipeline scheduler).
Figure 16: Execution timeline of the Pipeline-orchestration architecture.
Figure 17: The detailed illustration of the Rolling KV Cache and Sink Token designs.
Figure 18: Acceleration rate of different approaches on various resolutions. Left: Testing the acceleration rate of DiT only. Right: Testing the acceleration of the whole pipeline (with VAE).

### v2

@@ 3 Motivation & Bottleneck Analysis
In live-streaming scenarios, strict per-frame deadlines hinder the scalability of conventional parallelization strategies for two key reasons: (i) communication latency in sequence parallelism significantly reduces potential speedup, and (ii) short-frame chunks drive the workload into a memory-bound regime. These effects are further amplified in real-time streaming, where efficient causal DiTs operate on short sequences (e.g., 4 frames per step), reducing per-frame computation and making communication overhead proportionally heavier (see Fig. 4). We provide more information about the hardware intensity analysis and the configurations in Appendix Sec. 8.
@@ 5.2 Efficiency Evaluation
Figure 10: (a) The throughput results of different parallel inference systems on H100 GPUs. (b) The theoretical throughput of the proposed system and experimental results.
@@ 5.2.2 FPS Results
5.2.3 Throughput Analysis
Baseline throughput.
We report throughput results of Deepspeed-Ulysses jacobs2023deepspeed and Ring-Attention liu2023ring for comparison. Due to the communication overhead introduced by sequence parallelism, neither of these methods has achieved the throughput gain. A detailed result of the throughput gain is provided in Fig. 22, Appendix 10.
Theoretical throughput.
Following our performance-regime analysis and based on profiling, we conservatively assume a 50% rate bubble, and use the isolated VAE runtime $t_{\text{VAE}}$ as the ideal VAE inference time. The theoretical latency $t_{\text{theory}}$ is calculated as:
$t_{\text{theory}}=\frac{B/\text{bandwidth}}{\text{rate\_bubble}}+t_{\text{VAE}}.$ (3)
The theoretical throughput is given by
$\text{FPS}_{\text{theory}}=\frac{\text{chunk\_size}}{t_{\text{theory}}},$ (4)
and the results are shown in Fig. 10(b). The experimentally measured throughput closely follows the same scaling trend as the theoretical throughput, and approaches the theoretical values under certain configurations.
5.2.4 SLO-oriented Metrics
Table 1: Comparison of Latency, Jitter, and SLO Miss Rate between StreamDiffusionV2 and Baseline (CausVid with StreamVAE) on an H100 Workstation.
Method Metric (ms) Mean Std. P50 P90 P95 P99 Miss Rate (1s)
StreamDiffusionV2 (Ours) Tail Latency 357 361 484 497 585 0.2%
Jitter 21 30 13 45 61 132
CausVid Tail Latency 1760 1672 2958 3304 3896 99.9%
Jitter 235 255 164 498 599 1310
We report SLO-style tail latency together with jitter and miss rate metrics. Following common industry practice for interactive/live streaming [^4] https://www.mux.com/blog/the-low-latency-live-streaming-landscape-in-2019?, we target an end-to-end latency budget of 1.0s.
We provide the end-to-end latency distribution charts for our method and the baseline. As shown in Fig. 12, the baseline model, CausVid, has a dispersed distribution, and most samples have missed the 1s SLO deadline. These will also lead to severe jitter and lag during real-time applications. In contrast, our StreamDiffusionV2 achieves a concentrated distribution to meet the SLO.
For quantitative results, as shown in Tab. 1, on an H100 workstation, our method provides clear advantages in latency, stability, and SLO miss rate compared to the baseline. In particular, under a 1s SLO, the miss rate is only 0.2%, indicating that StreamDiffusionV2 reliably meets the latency budget. For stability, the jitter is 21 ms (mean) with 30 ms standard deviation, suggesting stable frame delivery under steady-state serving. All results are obtained from the online video-to-video experiments, conducted with a single-step configuration at a resolution of 512×512 on an H100 GPU.
@@ 5.3 Generation Quality Evaluation
Text-Image CLIP $\uparrow$ 26.48 - 27.69 29.29
Temporal CLIP $\uparrow$ 95.24 96.58 98.48 98.51
@@ 5.4.2 Effectiveness of the Dynamic DiT-Block Scheduler
To isolate algorithmic scaling, we evaluate the theoretical latency of sequence parallelism and pipeline parallelism with communication removed (Fig. 15). Our method, built on pipeline parallelism, attains a near-ideal acceleration by partitioning DiT blocks.
In contrast, sequence parallelism shows clear gains only at high resolutions; at moderate and low resolutions, it shifts the workload into a memory-bound regime, yielding little to no latency reduction and underutilizing compute.
@@ 6 Conclusions and New Outlooks
We propose StreamDiffusionV2, which closes the gap between offline video diffusion and live streaming constrained in real-time with SLO constraints. Our training-free system couples an SLO-aware batching/block scheduler with a sink-token–guided rolling KV cache, a motion-aware noise controller, and a pipeline orchestration that parallelizes across denoising steps and network stages—delivering near-linear FPS scaling without violating latency. It runs on heterogeneous GPUs and flexible step counts, achieving 0.5 s TTFF and up to 58.28 FPS (14B) / 64.52 FPS (1.3B) on 4×H100, and maintaining high FPS even as steps increase. These results make state-of-the-art generative live streaming practical for both individual creators and enterprise platforms.
Figure 17: The recent trends in the evolution of computational capability and memory bandwidth of AI hardware.
Figure 18: Intensity of the auto-regressive video diffusion model when the input sequence length varies.
Our new outlook. Video diffusion models have evolved from heavy bidirectional attention architectures toward autoregressive formulations, transforming a globally coupled spatiotemporal optimization into an amortized computation graph executed sequentially. From both hardware and algorithmic perspectives, this transition fundamentally reshapes the performance regime of video generation systems. We argue that future video models will increasingly operate in the memory-bound region.
  • Hardware trend. Across recent GPU generations (Fig. 18), from V100 to A100, H100, and GB100-class accelerators, compute capability has scaled significantly faster than memory bandwidth. For example, moving from V100 (Tensor Core peak $\approx 130$ TFLOP/s) to GB100 (dense FP16/BF16 Tensor Core peak $\approx 2.3$ PFLOP/s), peak compute increases by $\sim 17\times$, whereas HBM bandwidth grows from $\sim 1.1$ TB/s to $\sim 8$ TB/s (only $\sim 7\times$).
    Under the roofline model, this imbalance shifts the “knee” toward higher arithmetic intensity. Consequently, streaming inference workloads characterized by frequent memory movement and limited data reuse are increasingly constrained by memory bandwidth rather than raw compute. Therefore, the assumption that our serving pipeline is memory-sensitive is not transient; it becomes more pronounced as hardware continues to evolve.
  • Algorithm trend. On the algorithmic side, modern video generation models are progressively adopting more compressed and structured latent representations (e.g., stronger VAEs and tokenization schemes) to enable longer temporal horizons and larger spatiotemporal contexts under fixed compute and memory budgets.
    While compression reduces token counts and often improves reconstruction quality (e.g., Wan2.2 VAE reduces token counts to $\frac{1}{4}$ of Wan2.1), it also reduces arithmetic intensity by shrinking per-token compute relative to the required memory traffic in streaming inference. Our analysis shows that such compression can push the normalized arithmetic intensity to $\sim 0.2$ under the Wan2.2 VAE configuration (Fig. 18), further reinforcing the memory-bound regime in practical serving deployments.
Since both trends push the system deeper into the memory-bound regime, we believe that our approach is well-positioned to remain effective and potentially become more beneficial in future streaming systems by explicitly shaping memory traffic and scheduling under SLO constraints.